# Bayesian Uncertainty Quantification and Reliability Assessment for Mars Sample Return

### v1

@@ Abstract
In this paper, we employ a Bayesian approach to assess the reliability of a critical component in the Mars Sample Return program, focusing on the Earth Entry System’s risk of containment not assured upon reentry. Our study uses Gaussian Process modeling under a Bayesian regime to analyze the Earth Entry System’s resilience against operational stress. This Bayesian framework allows for a detailed probabilistic evaluation of the risk of containment not assured, indicating the feasibility of meeting the mission’s stringent safety goal of $0.999999$ probability of success. The findings underscore the effectiveness of Bayesian methods for complex uncertainty quantification analyses of computer simulations, providing valuable insights for computational reliability analysis in a risk-averse setting.
@@
1 Introduction
In July 2020 the National Aeronautics and Space Administration (NASA) launched the Mars Perseverance Rover (Farley et al.,, 2020). Perseverance landed on Mars in February 2021 with the main goal of seeking signs of ancient life by the collection of rock cores, regolith and atmospheric samples for later evaluation on Earth (Beaty et al.,, 2019; Haltigin et al.,, 2022; Grady et al.,, 2022; Swindle et al.,, 2022; Carrier et al.,, 2022). Over the past several years, as Perseverance has been collecting samples, NASA and the European Space Agency (ESA) have been formulating the next elements of the Mars Sample Return (MSR) campaign. Prior to the re-architecture efforts undertaken in response to the second MSR Independent Review Board report (Figueroa et al.,, 2023), MSR would gather the sealed sample tubes, place them inside the Orbiting Sample (OS) container, launch them into Mars orbit, capture and contain them securely, and returm them safely to Earth through an Earth Entry System (EES) (Sarli et al.,, 2024).
In line with current policies (United Nations General Assembly,, 1967; COSPAR Panel on Planetary Protection,, 2021), the design, planning and execution of MSR must meet planetary protection requirements (NASA Office of Safety and Mission Assurance,, 2021, 2022). In this manuscript, we focus on Backward Planetary Protection (BPP), that is, protecting the Earth-Moon system from possible harmful extraterrestrial contamination that may be returned from Mars. These requirements translate into establishing and implementing a strategy and design concepts to break the chain of contact with Mars by isolating and robustly containing the restricted samples (Cataldo et al.,, 2024). To achieve compliance with these BPP requirements, quantitative targets have been established that correspond to a $0.999999$ probability of maintaining containment. Achieving this requires a statistical model involving both reliability analysis and extreme value distributions to assess the probability of a rare event, that is, the one-in-a-million probability of containment not assured, $\text{P}_{CNA}$. To evaluate the feasibility of this probabilistic goal, we focus on the assessment of the landing performance. In so doing, this paper specifically informs the process and prospects for complying with BPP policies using an EES-like vehicle.
In this analysis, we explore a particular cause of containment not assured when the maximum peak acceleration of the OS upon landing exceeds 3000 Gs. In particular, we examine the effects of uncertainty in specific materials properties of individual EES components. We note that while materials might react to additional unknowns affecting their behavior, this is beyond the scope of this paper. The uncertainty being used here is only due to the inherent variability of these materials properties. Assessing the reliability and $\text{P}_{CNA}$ of such complex systems involves significant uncertainty. Traditional deterministic methods may not be adequate to address such uncertainties due to inherent randomness and model parameter uncertainty. The Bayesian methodology, as a probabilistic approach, has been increasingly recognized as a powerful tool for managing uncertainty and variability in reliability analysis (Sankararaman and Mahadevan,, 2011, 2015). The aim of this study is to develop Bayesian methods that allow for model and parameter uncertainty as well as the inherent randomness of the system when assessing the probability of loss of containment.
@@ 1 Introduction
In the following section, an overview of the data utilized is provided. In Section 3, we outline the methodology applied during our analysis. Section 4 is devoted to the application of these methods with respect to our specific data set as well as the results of the analysis. Section 5 provides a conclusion and recommendation for future work.
2 Data
@@
For the use-case investigated here, earlier work completed by the Jet Propulsion Laboratory (JPL) identified 15 input variables comprising Young’s Modulus, Shear Modulus, and compressive and tensile strength in varying directions for two different materials, IM7 (a carbon fiber) and Kevlar (Carpenter,, 2021). Experimental tests executed by JPL provided initial observations for each of these 15 input variables. The number of observations for each variable ranges from 3 to 12. Previous Uncertainty Quantification (UQ) and reliability analysis work (Naresh et al.,, 2018; Fitt et al.,, 2019) as well as typical engineering practice suggest fitting two-parameter Weibull distributions to the strength variables and Normal distributions to the Modulus variables. Because of this precedence, as well as our small sample sizes and the need for extrapolation, we assume Weibull and Normal distributions, respectively. We use the indicators $X0001-X0015$ to identify the input variables moving forward and refer the reader to the supplementary materials for more details regarding the data.
@@
After determining the parameters of the respective Normal and Weibull distributions for the 15 input variables using Maximum Likelihood Estimation (MLE), JPL engineers generated 25 samples of the input variables using Latin Hypercube Sampling (LHS). Using these 25 samples, the JPL engineers then ran 25 simulations of an LS-DYNA EES System-Level Impact Model (SLIM) and extracted peak OS acceleration as the quantity of interest (Livermore Software Technology, An Ansys Company,, 2021; Siddens et al.,, 2022, 2023). In the context of our analysis, we consider the 25 realization of peak OS acceleration as our output variable of interest. With the particulars of the data in mind, we next review the methods applied throughout our analysis.
3 Methods
@@
In the following section, we outline the various methodologies employed in our analysis. A direct MLE approach is not appropriate because of the high dimension of the variable space and the low sample size; though we compare a functional regularized Restricted Maximum Likelihood (REML) approach to our preferred Bayesian method, which we favor due to its effective ability to incorporate uncertainty in the parameters. We begin with a consideration of the priors applied within the Bayesian analysis of our input variables.
3.1 Prior considerations
@@
We have 15 input variables, each represented by either a Weibull or a Normal distribution; our goal with the Bayesian approach is to generate posterior distributions of the input variables’ parameters. For all the input variables, regardless of distributional assumptions, we compare the frequentist estimates provided by the MLE values and the mean of the posterior distributions generated by several priors within a Bayesian framework. That is, we assume a prior density on the unknown parameters $\pi_{0}(\psi)$ (here $\psi=(\alpha,\beta)$ or $\psi=(\mu,\sigma^{2})$ according to distributional assumption). Then, by applying Bayes’ Rule, we have $\pi(\psi\mid X_{1},\ldots,X_{n})=\frac{\pi_{0}(\psi)\prod_{i=1}^{n}{f(X_{i}\mid\psi)}}{\int\pi_{0}(\psi^{\prime})\prod_{i=1}^{n}{f(X_{i}\mid\psi^{\prime})}d\psi^{\prime}}$, where the distribution $\pi(\psi\mid X_{1},\ldots,X_{n})$ is known as the posterior distribution and reflects the updated knowledge of the parameters conditional on the data (Gelman et al.,, 2013).
@@
3.2 Adaptive Metropolis (AM) algorithm
@@
3.3 Gaussian processes and spatial statistics
@@
First we consider the basic approach to modeling complex processes using computer simulations. Given a selection of input variables $\mathbf{X}=(X_{1},...,X_{k})$, an output variable, $Z=h(\mathbf{X})$, is produced by the computer code. In many instances, the goal of generating the output variable is prediction. Sacks et al., (1989) discusses the use of stochastic processes in modeling the prediction of the response variable while Kennedy and O’Hagan, (2001) expand upon this by using Bayesian methods in the calibration of computer models to improve the prediction process.
@@
For our study, we model the output, $Z$, using a Gaussian Process (GP), a widely-used tool for dealing with spatially structured data (Rasmussen and Williams,, 2006). The GP provides a flexible, nonparametric approach, which is particularly apt at capturing spatial correlations (Cressie,, 1993). We define the GP model as $Z\sim N_{n}(X\beta,\Sigma)$, where Z represents the vector of simulation outputs, $X\beta$ denotes the mean response as a function of parameters $\beta$ ($X$ represents the design matrix of the GP, not the set of input variables previously defined as $\mathbf{X}$), and $\Sigma$ is the covariance matrix capturing the dependence structure among the data.
@@
In GP s, the covariance between any two observations depends on the input locations corresponding to these observations. That is, we let $\Sigma=\alpha V(\mathbf{\theta})$, where $\alpha>0$ is a scale parameter and $V$ is a function of the spatial range parameters $\mathbf{\theta}$. Consistent with previous research (Sacks et al.,, 1989), $V$ is the squared exponential function. Assuming this form allows for an anisotropic covariance structure by using different spatial range parameters $\rho_{k}$, which scale the distances along each input dimension $k$, thus capturing varying influences of the inputs. We define $V$ by:
$v_{ij}=\exp{\bigg{(}-\sum_{k=1}^{K}\frac{(s_{ik}-s_{jk})^{2}}{\rho_{k}^{2}}\bigg{)}}$ (1)
where $s_{ik}$ and $s_{jk}$ are the $k$-th input variables of observations $i$ and $j$. We ensure the necessary positivity of the spatial range parameters for the Gaussian process model by setting $\theta_{k}=\log{\rho_{k}}$ (Rasmussen and Williams,, 2006).
To estimate our spatial range parameters, we again apply a Bayesian framework and compare those results to the MLE values. We consider several options using varying likelihood equations as outlined by Stein, (1999). We include some relevant equations as well as the likelihood equations in Table 1 but refer the interested reader to Stein’s book for further reference (Stein,, 1999).
Within our analysis we will employ the regularized REML Negative Log Likelihood (NLLH):
$l_{R}(\theta)=l_{w}(\theta)+\lambda\sum_{k=1}^{K}(\theta_{k}-\bar{\theta})^{2}$ (2)
for our MLE results. This particular formulation was incorporated in order to stabilize the estimation process which encountered issues due to the relatively small size of our data set. For the Bayesian portion of our analysis, we use the Bayesian NLLH:
@@
where $\pi(\theta)$ is the prior distribution (Cressie,, 1993).
probability density function (pdf) $f(Z|X\beta,\Sigma)=(2\pi)^{-\frac{n}{2}}|\Sigma|^{-\frac{1}{2}}\exp\left(-\frac{1}{2}[Z-X\beta]^{T}\Sigma^{-1}[Z-X\beta]\right)$
NLLH $l(\beta,\alpha,\theta)=\frac{n}{2}\log(2\pi)+\frac{n}{2}\log(\alpha)+\frac{1}{2}\log{|V(\theta)|}+\frac{1}{2\alpha}[Z-X\beta]^{T}V(\theta)^{-1}[Z-X\beta]$
Estimates $\hat{\beta}(\theta)=(X^{T}V(\theta)^{-1}X)^{-1}X^{T}V(\theta)^{-1}Z,\quad\hat{\alpha}(\theta)=\frac{G^{2}(\theta)}{n}$
Sub-expressions $H=V(\theta)^{-1}-V(\theta)^{-1}X(X^{T}V(\theta)^{-1}X)^{-1}X^{T}V(\theta)^{-1},\quad G^{2}(\theta)=Z^{T}HZ$
Profile NLLH $l_{p}(\theta)=l(\hat{\beta}(\theta),\hat{\alpha}(\theta),\theta)=\frac{n}{2}\log(2\pi)+\frac{n}{2}\log\left(\frac{G^{2}(\theta)}{n}\right)+\frac{1}{2}\log{|V(\theta)|}+\frac{n}{2}$
REML NLLH $l_{w}(\theta)=\frac{n-q}{2}\log(2\pi)+\frac{n-q}{2}\log\left(\frac{G^{2}(\theta)}{n-q}\right)-\frac{1}{2}\log{|X^{T}X|}+\frac{1}{2}\log{|X^{T}V(\theta)^{-1}X|}+\frac{1}{2}\log{|V(\theta)|}+\frac{n-q}{2}$
Table 1: NLLH Equations
In the next section, we discuss the use of these varying methods, to include the Bayesian method with the prior $\theta_{1:K}\sim N(\tau,\nu^{2})$. We use Leave One Out Cross Validation (CV) twice, first to make an appropriate choice for the penalty term in the regularized REML case and again when choosing the hyperparameters $(\tau,\nu^{2})$ in the Bayesian case.
Once we have generated the posterior distributions for $\theta_{1:K}$, we can then predict new values from the computer simulations using a kriging approach. Kriging, a geostatistical interpolation technique, provides optimal, unbiased predictions by utilizing the spatial correlation structure inferred by the Gaussian process model (Cressie,, 1993; Stein,, 1999). We apply ordinary kriging to the problem, aiming to predict a new value $z_{0}$ with mean $x_{0}^{T}\beta$, variance $\sigma_{0}^{2}$, and covariance $\text{Cov}(z_{0},\mathbf{Z})=\phi$. We consider predictors of the form $\hat{z}_{0}=\gamma^{T}Z$ subject to the constraint $\gamma^{T}X=x_{0}^{T}$. Following the Lagrange Multiplier approach, we end up minimizing the mean squared prediction error (MSPE): $E[(z_{0}-\hat{z_{0}})^{2}]=\sigma_{0}^{2}-2\gamma^{T}\phi+\gamma^{T}\Sigma\gamma$ subject to the above constraint. Thus, our predictor and MSPE are:
@@
After we have generated the posterior distributions for the parameters of the Normal and Weibull variables as well as for the spatial range parameters, we are able to predict the peak OS acceleration. A distinct advantage of the Bayesian approach to generate the spatial range parameters, as opposed to estimating the spatial process using MLE values for $\theta_{1:K}$, is that we are accounting for the error in the estimation of the spatial range parameters. This provides a more robust quantification of the underlying uncertainty within the model. With these elements in place, we can then simulate the posterior distribution of the probability of exceeding 3000 Gs with the given information.
3.4 Cross validation in a Bayesian setting
We mentioned in Section 3.3 that we would be using CV in two instances. In the first scenario, when choosing the penalty term for the regularized REML case, our CV follows the methods outlined by Hastie et al., (2009). For both uses of CV, we employ a squared error loss function. The algorithm used for the first CV is outlined in Algorithm 1. The input values for the algorithm are the input variables, $\mathbf{X}_{n\times K}$, the output variable, $\mathbf{Z}_{n\times 1}$, and the selected potential values for $\lambda$, $\mathbf{\Lambda}_{Q\times 1}$. The output of the algorithm is the vector of CV scores for each value of $\lambda$ used, $\mathbf{L}_{Q\times 1}$.

### v2

# A Case Study on Quantifying Reliability under Extreme Risk Constraints in Space Missions

@@ Abstract
In this paper, we employ a Bayesian approach to uncertainty quantification of computer simulations used to assess the probability of rare events. As a case study, we assess the reliability of an Earth reentry capsule for sample return missions that must be able to withstand the reentry loads in order to land intact. Our study uses Gaussian Process modeling under a Bayesian regime to analyze the reentry vehicle’s resilience against operational stress. This Bayesian framework allows for a detailed probabilistic evaluation of the system’s reliability, indicating our ability to verify stringent safety goals of rare events with a $0.999999$ of probability of success. The findings underscore the effectiveness of Bayesian methods for complex uncertainty quantification analyses of computer simulations, providing valuable insights for computational reliability analysis in a risk-averse setting.
@@
1 Problem description
In reliability analysis, quantifying the probability of a rare event, such as a system failure occurring with probability less than one in a million, poses a significant challenge. When the system is evaluated via an expensive computer simulation, and when only limited data are available to characterize input uncertainty, traditional methods of uncertainty quantification (UQ) may not be adequate. This issue is further complicated in settings where failure is defined in terms of exceeding an extreme threshold, meaning that both interpolation and extrapolation become critical.
In such cases, surrogate modeling is typically used to emulate the computer code and provide predictions at untested input values. Gaussian Process (GP) modeling has become a standard approach in this context, offering both flexibility and interpretability (Sacks et al.,, 1989; Stein,, 1999; Rasmussen and Williams,, 2006). These surrogate models are often paired with simulation-based techniques to estimate quantities of interest, such as failure probabilities, under a probabilistic framework. However, when input data are sparse, the uncertainty in the input distributions themselves must be incorporated into the analysis. Failing to do so can produce misleading estimates, particularly when the rare event lies in the tail of the response distribution.
Bayesian methods provide a natural framework for incorporating both parameter and model uncertainty in this setting. Prior work has established Bayesian approaches to reliability analysis (Sankararaman and Mahadevan,, 2011, 2015) and Bayesian calibration of computer models (Kennedy and O’Hagan,, 2001; O’Hagan,, 2006; Bayarri et al.,, 2007). These techniques allow for formal updating of uncertain model parameters and for coherent propagation of uncertainty through the simulation process. More recent work has also addressed practical challenges in surrogate modeling for complex systems, including computational efficiency, heteroskedasticity, and high-dimensional input spaces (Gramacy,, 2020; Binois et al.,, 2018).
Despite these developments, relatively few studies have addressed the combination of three factors critical in extreme reliability assessment: (1) sparse or limited empirical data on input distributions; (2) surrogate modeling of simulator outputs via GPs, where the spatial correlation structure itself is uncertain; and (3) estimation of very low tail probabilities via posterior simulation. In particular, while GP modeling is well established, the spatial range parameters are often fixed via maximum likelihood, rather than treated probabilistically and integrated into the overall uncertainty. The impact of this modeling choice can be especially pronounced when predicting values in the tails, where no simulator runs exist and where extrapolation is necessary.
This paper presents a fully Bayesian approach for estimating rare event probabilities in simulation-based reliability analysis. The input variables are assumed to follow known distributional forms, with parameters estimated from small-sample data using Bayesian methods. The output of a deterministic computer simulation is modeled using a Gaussian process (GP), with prior distributions placed on the spatial range parameters to account for correlation uncertainty in the surrogate. Cross-validation is used to tune hyperparameters and regularization penalties. Posterior distributions for all model components are then propagated through to the output prediction, and the final quantity of interest, the probability that the simulator output exceeds a critical threshold (which we call probability of failure $(\text{P}_{f})$), is estimated via posterior simulation. A safety-critical system with one-in-a-million reliability requirements is used to illustrate the method.
@@ 1 Problem description
In the following section, an overview of the data utilized is provided. In Section 3, we outline the methodology applied during our analysis, as well as the results of the analysis in this case study. Section 4 provides a recommendations for employing the results of our work in practice and suggestions for future work based on our findings.
2 Data collection and preparation
The data used in this study originate from a set of structural simulations intended to evaluate the preliminary design response of the Earth reentry system of a space sample return mission under high-impact loading (Cataldo et al., 2025a, ). Engineers used a finite element model to simulate the capsule’s Earth reentry. The simulations focused on how variation in material properties, specifically strength and stiffness parameters, would affect the peak acceleration of the sample container experienced under nominal conditions. Empirical measurements of these properties were obtained through laboratory testing of two composite materials (in varying directions), and were then used to define the uncertain input variables for the model.
@@ 2 Data collection and preparation
More specifically, engineers identified 15 input variables comprising Young’s modulus, shear modulus, and compressive and tensile strength in varying directions for two different materials, IM7 (a carbon fiber) and Kevlar (Carpenter,, 2021). Experimental tests executed by the engineers provided initial observations for each of these 15 input variables. The number of observations for each variable ranges from 3 to 12. Previous Uncertainty Quantification (UQ) and reliability analysis work (Naresh et al.,, 2018; Fitt et al.,, 2019) as well as typical engineering practice suggest fitting two-parameter Weibull distributions to the strength variables and Normal distributions to the Modulus variables. Because of this precedence, as well as our small sample sizes and the need for extrapolation, we assume Weibull and Normal distributions, respectively. We use the indicators $X0001-X0015$ to identify the input variables moving forward as can be seen in Table 1.
X0001 $IM7-E_{a},E_{b},E_{c}$ Young’s Modulus 8 Normal
X0002 $IM7-G_{ab},G_{bc},G_{ca}$ Shear Modulus 4 Normal
X0003 $IM7-X_{c}$ Compressive Strength (0 Deg) 3 Weibull
X0004 $IM7-X_{t}$ Tensile Strength (0 Deg) 4 Weibull
X0005 $IM7-Y_{c}$ Compressive Strength (90 Deg) 3 Weibull
X0006 $IM7-Y_{t}$ Tensile Strength (90 Deg) 4 Weibull
X0007 $IM7-S_{c}$ Shear Strength 4 Weibull
X0008 $Kv-E_{a},E_{c}$ Young’s Modulus (0 Deg) 4 Normal
X0009 $Kv-E_{b}$ Young’s Modulus (90 Deg) 4 Normal
X0010 $Kv-G_{ab},G_{bc},G_{ca}$ Shear Modulus 4 Normal
X0011 $Kv-X_{c}$ Compressive Strength (0 Deg) 12 Weibull
X0012 $Kv-X_{t}$ Tensile Strength (0 Deg) 4 Weibull
X0013 $Kv-Y_{c}$ Compressive Strength (90 Deg) 11 Weibull
X0014 $Kv-Y_{t}$ Tensile Strength (90 Deg) 4 Weibull
X0015 $Kv-S_{c}$ Shear Strength 4 Weibull
Table 1: Description of input variables
@@
After determining the parameters of the respective Normal and Weibull distributions for the 15 input variables using Maximum Likelihood Estimation (MLE), the engineers next generated 25 samples of the input variables using Latin Hypercube Sampling (LHS). Using these 25 samples, the engineers then ran 25 simulations of an LS-DYNA Earth entry vehicle’s system-level impact model and extracted the peak acceleration as the quantity of interest (Livermore Software Technology, An Ansys Company,, 2021; Siddens et al.,, 2022, 2023). In the context of our analysis, we consider the 25 realization of the peak acceleration as our output variable of interest. With the particulars of the data in mind, we next review the methods applied throughout our analysis.
3 Analysis and interpretation
3.1   Methods
@@
In the following sections, we outline the various methodologies employed in our analysis. A direct MLE approach is not appropriate because of the high dimension of the variable space and the low sample size; though we compare a functional regularized Restricted Maximum Likelihood (REML) approach to our preferred Bayesian method, which we favor due to its effective ability to incorporate uncertainty in the parameters. We begin with a consideration of the priors applied within the Bayesian analysis of our input variables.
3.1.1   Prior considerations
@@
We have 15 input variables, each represented by either a Weibull or a Normal distribution; our goal with the Bayesian approach is to generate posterior distributions of the input variables’ parameters. For all the input variables, regardless of distributional assumptions, we compare the frequentist estimates provided by the REML values and the mean of the posterior distributions generated by several priors within a Bayesian framework. That is, we assume a prior density on the unknown parameters $\pi_{0}(\psi)$ (here $\psi=(\alpha,\beta)$ for Weibull variables or $\psi=(\mu,\sigma^{2})$ for Normal variables). Then, by applying Bayes’ Rule, we have $\pi(\psi\mid X_{1},\ldots,X_{n})=\frac{\pi_{0}(\psi)\prod_{i=1}^{n}{f(X_{i}\mid\psi)}}{\int\pi_{0}(\psi^{\prime})\prod_{i=1}^{n}{f(X_{i}\mid\psi^{\prime})}d\psi^{\prime}}$, where the distribution $\pi(\psi\mid X_{1},\ldots,X_{n})$ is known as the posterior distribution and reflects the updated knowledge of the parameters conditional on the data (Gelman et al.,, 2013).
@@
3.1.2   Adaptive Metropolis (AM) algorithm
@@
3.1.3   Gaussian processes and spatial statistics
@@
First we consider the basic approach to modeling complex processes using computer simulations. Given a selection of input variables $\mathbf{X}=(X_{1},...,X_{k})$, an output variable, $Z=h(\mathbf{X})$, is produced by the computer code. In many instances, the reason for generating such output variables is to then use the results for making further predictions. Sacks et al., (1989) discusses the use of stochastic processes in modeling the response and making predictions while Gramacy, (2020) expand upon this by using Bayesian methods in the calibration of computer models to improve the prediction process.
@@
For our study, we model the output, $Z$, using a GP, a widely-used tool for dealing with spatially structured data (Rasmussen and Williams,, 2006). The GP provides a flexible, nonparametric approach, which is particularly apt at capturing spatial correlations (Cressie,, 1993). We define the GP model as $Z\sim N_{n}(X\beta,\Sigma)$, where Z represents the vector of simulation outputs, $X\beta$ denotes the mean response as a function of parameters $\beta$ ($X$ represents the design matrix of the GP, not the set of input variables previously defined as $\mathbf{X}$), and $\Sigma$ is the covariance matrix capturing the dependence structure among the data.
@@
In GP s, the covariance between any two observations depends on the input locations corresponding to these observations. That is, we let $\Sigma=\alpha V(\mathbf{\theta})$, where $\alpha>0$ is a scale parameter and $V$ is a function of the spatial range parameters $\mathbf{\theta}$ and the euclidean distances between observations. Considered a popular choice for such applications, we use the squared exponential function for $V$ (Rasmussen and Williams,, 2006; Hadji and Szabó,, 2019). This form allows for an anisotropic covariance structure by using different spatial range parameters $\theta_{k}$, which scale the distances along each input dimension $k$, thus capturing varying influences of the inputs. We define $V$ by its individual entries:
$v_{ij}=\exp{\bigg{(}-\sum_{k=1}^{K}\frac{(x_{ik}-x_{jk})^{2}}{\exp(\theta_{k})^{2}}\bigg{)}}$ (1)
where $x_{ik}$ and $x_{jk}$ are the $k$-th input variables of observations $i$ and $j$.
In order to estimate our spatial range parameters, we will again apply a Bayesian framework; however, we begin with outlining the MLE approach as presented by Stein (1999). Given our definition of $Z$ as above, we have:
$f(Z|X\beta,\Sigma)=(2\pi)^{-n/2}|\Sigma|^{-1/2}\exp{\bigg{(}-\frac{1}{2}[Z-X\beta]^{T}\Sigma^{-1}[Z-X\beta]\bigg{)}}$ (2)
This produces the negative log likelihood:
$l(\beta,\alpha,\theta)=\frac{n}{2}\log{(2\pi)}+\frac{n}{2}\log{\alpha}+\frac{1}{2}\log{|V(\theta)|}+\frac{1}{2\alpha}[Z-X\beta]^{T}V(\theta)^{-1}[Z-X\beta]$
Using the least squares estimator of $\beta$ based on the covariance matrix $V(\theta)$:
$\hat{\beta}(\theta)=(X^{T}V(\theta)^{-1}X)^{-1}X^{T}V(\theta)^{-1}Z$ (3)
and defining:
$H=V(\theta)^{-1}-V(\theta)^{-1}X(X^{T}V(\theta)^{-1}X)^{-1}X^{T}V(\theta)^{-1}$ and $G^{2}(\theta)=Z^{T}HZ$, we then have:
$l(\hat{\beta}(\theta),\alpha,\theta)=\frac{n}{2}\log{(2\pi)}+\frac{n}{2}\log{\alpha}+\frac{1}{2}\log{|V(\theta)|}+\frac{1}{2\alpha}G^{2}(\theta)$
Finally, we can minimize this analytically with respect to $\alpha$, resulting in: $\hat{\alpha}(\theta)=\frac{G^{2}(\theta)}{n}$, giving us:
$l^{*}(\theta)=l(\hat{\beta}(\theta),\hat{\alpha}(\theta),\theta)=\frac{n}{2}\log{(2\pi)}+\frac{n}{2}\log{\frac{G^{2}(\theta)}{n}}+\frac{1}{2}\log{|V(\theta)|}+\frac{n}{2}$
The above quantity is often called the profile negative log likelihood.
Another option in terms of likelihoods that arises in the realm of spatial statistics is the restricted maximum likelihood or REML approach, which was first introduced by Patterson and Thompson (1971). The idea behind the REML method is to separate the two part estimation problem, that of estimating the linear model and the estimation of the covariance structure. This is done by considering the likelihood function of the contrasts; that is, let $W=A^{T}Z$ be a vector of $n-q$ linearly independent contrasts, then $W\sim N(0,A^{T}\Sigma A)$. This gives the pdf:
$f(W\mid A,\Sigma)=(2\pi)^{-(n-q)/2}|A^{T}\Sigma A|^{-1/2}\exp{\left(-\frac{1}{2}W^{T}|A^{t}\Sigma A|^{-1}W\right)}$
with the negative log likelihood:
$l_{w}(\alpha,\theta)=\big{(}\frac{n-q}{2}\big{)}\log{(2\pi\alpha)}+\frac{1}{2}\log{|A^{T}V(\theta)A|}+\frac{1}{2}W^{T}|A^{T}V(\theta)A|^{-1}W$
Following the calculations of Patterson and Thompson (1971) and suggestions made by Harville (1977), the above likelihood can be simplified to:
$l_{w}(\alpha,\theta)=\big{(}\frac{n-q}{2}\big{)}\log{(2\pi\alpha)}-\frac{1}{2}\log{X^{T}X}+\frac{1}{2}\log{|X^{T}V(\theta)^{-1}X|}+\frac{1}{2}\log{|V(\theta)|}+\frac{1}{2\alpha}G^{2}(\theta)$
Finally, minimizing with respect to $\alpha$, results in $\tilde{\alpha}=\frac{G^{2}(\theta)}{n-q}$, giving us:
$\begin{split}l_{w}^{*}(\theta)=\frac{n-q}{2}\log{2\pi}+\frac{n-q}{2}\log{\frac{G^{2}(\theta)}{n-q}}&-\frac{1}{2}\log{|X^{T}X|}\\ &+\frac{1}{2}\log{|X^{T}V(\theta)^{-1}X|}+\frac{1}{2}\log{|V(\theta)|}+\frac{n-q}{2}\end{split}$ (4)
Therefore, $l_{w}^{*}(\theta)$ is what we will refer to as the REML negative log likelihood.
Before moving on to the Bayesian approach, we include a brief mention of a regularized REML method. With some models, there might exist stabilization issues with the estimates; or, estimation problems can arise when the number of parameters exceeds the number of data points. The regularized REML approach works to prevent these issues by including a penalty term to the likelihood function. We will apply a Ridge-type regularization where our regularized REML negative log-likelihood takes the form:
$l_{R}(\theta)=l_{w}^{*}(\theta)+\lambda\sum_{k=1}^{K}(\theta_{k}-\bar{\theta})^{2}$ (5)
Here, the penalty term is $\lambda$ and the L2 penalty function tends to shrink the spatial range parameters towards their mean value.
Finally, we turn to the Bayesian approach where a prior, $\frac{\pi(\theta)}{\alpha}$, is placed on the spatial range parameters. Using our defined pdf from Equation 2 and defining $\hat{\beta}$ as in 3 we have the following posterior distribution:
$\pi(\beta,\alpha,\theta\mid Z)\propto\frac{\pi(\theta)}{\alpha}\alpha^{-n/2}|V(\theta)|^{-1/2}$ $\cdot\exp{\bigg{[}-\frac{G^{2}(\theta)}{2\alpha}\bigg{]}}$
$\cdot\exp{\bigg{[}-\frac{1}{2\alpha}(\beta-\hat{\beta})^{T}X^{T}V(\theta)^{-1}X(\beta-\hat{\beta})\bigg{]}}$
Integrating out with respect to $\beta$ and $\alpha$ we have:
$\pi(\theta\mid Z)\propto\pi(\theta)|V(\theta)|^{-1/2}G^{2}(\theta)^{\frac{q-n}{2}}|X^{T}V(\theta)^{-1}X|^{-1/2}$
Which gives us the Bayesian negative log likelihood of:
@@
While we presented the derivation of several versions of applicable negative log-likelihood equations, we note here that we utilize the regularized REML as the frequentist version in comparison to the Bayesian approach. In the next section, we discuss the use of these varying methods, to include the Bayesian method with the prior $\theta_{1:K}\sim N(\tau,\nu^{2})$. We use Leave One Out Cross Validation (CV) twice, first to make an appropriate choice for the penalty term in the regularized REML case and again when choosing the hyperparameters $(\tau,\nu^{2})$ in the Bayesian case.
Once we have generated the posterior distributions for $\theta_{1:K}$, we can then predict new values from the computer simulations using a kriging approach. Kriging, a geostatistical interpolation technique, provides optimal, unbiased predictions by utilizing the spatial correlation structure inferred by the Gaussian process model (Cressie,, 1993; Stein,, 1999). We apply ordinary kriging to the problem, aiming to predict a new value $z_{0}$ with mean $x_{0}^{T}\beta$, variance $\sigma_{0}^{2}$, and covariance $\text{Cov}(z_{0},\mathbf{Z})=\phi$. Following the Lagrange Multiplier approach with Lagrange multiplier $\gamma$, we consider predictors of the form $\hat{z}_{0}=\gamma^{T}Z$ subject to the constraint $\gamma^{T}X=x_{0}^{T}$. Thus, we end up minimizing the mean squared prediction error (MSPE): $E[(z_{0}-\hat{z_{0}})^{2}]=\sigma_{0}^{2}-2\gamma^{T}\phi+\gamma^{T}\Sigma\gamma$ subject to the above constraint. This gives us the predictor and MSPE:
@@
After we have generated the posterior distributions for the parameters of the Normal and Weibull variables as well as for the spatial range parameters, we are able to predict the peak acceleration of the sample container. A distinct advantage of the Bayesian approach to generate the spatial range parameters, as opposed to estimating the spatial process using REML values for $\theta_{1:K}$, is that we are accounting for the error in the estimation of the spatial range parameters. This provides a more robust quantification of the underlying uncertainty within the model. With these elements in place, we can then simulate the posterior distribution of the probability of exceeding 3000 Gs with the given information.
3.1.4   Cross validation in a Bayesian setting
We mentioned in Section 3.1.3 that we would be using CV in two instances. In the first scenario, when choosing the penalty term for the regularized REML case, our CV follows the methods outlined by Hastie et al., (2009). For both uses of CV, we employ a squared error loss function. The algorithm used for the first CV is outlined in Algorithm 1. The input values for the algorithm are the input variables, $\mathbf{X}_{n\times K}$, the output variable, $\mathbf{Z}_{n\times 1}$, and the selected potential values for $\lambda$, $\mathbf{\Lambda}_{Q\times 1}$. The output of the algorithm is the vector of CV scores for each value of $\lambda$ used, $\mathbf{L}_{Q\times 1}$.